OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). -1 1 0 1/2

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ -3}})\quad &({\color{red}{ 3}}\quad ,&{\color{blue}{ -5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (haseeb96):

y2-y1/x2-x1 =m use this equation and get the answer

OpenStudy (wolf1728):

slope = change in y / change in x slope = (-3 -(-5)) / (1-3) slope = (-3 +5) / (-2) slope = 2 / -2 slope = -1

OpenStudy (anonymous):

as @jdoe0001 said just simply plug in the numbers and subtract. Easy. (x1,y1) (x2,y2) (1, -3) (3, -5) \[ \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}=\]